OpenStudy (anonymous):

given that one zero is -4, find all zeros of p(x)=X^3+x^2-22x-40 a) the root of the equation are -5,-4, and -2 b) the root of the equation are 7,-4, and 3 c) the root of the equation are 10,-4 and 2 d) the root of the equation are 5, -4 and -2

OpenStudy (anonymous):

Assuming you know long division, divide and then factor.

OpenStudy (campbell_st):

well there is a short cut if x = -4 is a zero then (x + 4) is a factor and you know its a cubic so y = (x + a)(a+b)(x + 4) and the value of 4ab the product of the constants on the binomials is -40 so therefore 4ab = -40 or ab = -10 so either a or b is positive.... but not both match then to your answers y=(x -5)(x + 2)(x +4) or y = (x + 5)(x -2)(x+4) find the corresponding zeros from the list

OpenStudy (anonymous):

just do synthetic long division, (x+4) (x^2 - 3x - 10) (x+4) (x-5)(x+2) the other zeros are 5 and -2

OpenStudy (anonymous):

so it D??

OpenStudy (anonymous):

yes, it's D

OpenStudy (campbell_st):

you don't need any division if you know about the connect to the constant in the expanded form and yes D

OpenStudy (anonymous):

thanks alot guys!!